OpenStudy (lena772):

The equation of a line is shown below. y = fraction reading 1 over 5x - 9 What is the equation of a line which is perpendicular to this line and passes through (1, 58)? y = 5x - 63 y = -5x + 63 y = 9x - 63 y = 9x + 63

ganeshie8 (ganeshie8):

wats the slope of the GIVEN line ?

ganeshie8 (ganeshie8):

\(\large y = \frac{1}{5}x - 9\)

OpenStudy (lena772):

1/5

ganeshie8 (ganeshie8):

compare it wid \\(\large y = mx + b\)

ganeshie8 (ganeshie8):

Yes, slope of given line is 1/5

ganeshie8 (ganeshie8):

we want to find the slope of line thats perpendicular to this

ganeshie8 (ganeshie8):

u knw any relation between slopes, when lines are perpendicular ?

OpenStudy (lena772):

The recipricol right?

OpenStudy (lena772):

negative recipricol

ganeshie8 (ganeshie8):

half correct

ganeshie8 (ganeshie8):

yes! negative reciprocal !!

OpenStudy (lena772):

:)

ganeshie8 (ganeshie8):

so what will be the slope of the required line ?

OpenStudy (lena772):

So it's -1/(1/5)

ganeshie8 (ganeshie8):

simplify

OpenStudy (lena772):

-1/0.2 = -5?

ganeshie8 (ganeshie8):

yup ! you knw slope \(-5\), and u knw a point \((1, 58)\)

ganeshie8 (ganeshie8):

can u write the equaiton of line ?

OpenStudy (lena772):

y=-5x+b 58=-5x+b 58=-5+b 63=b?

ganeshie8 (ganeshie8):

Awesome :)

OpenStudy (lena772):

thanks

ganeshie8 (ganeshie8):

np :)